OpenStudy (anonymous):

1/3 minus -7/8 help? ((Algebra 1))

OpenStudy (anonymous):

is it double minus?

OpenStudy (stefrheart):

you need to make it common denominator

OpenStudy (anonymous):

29/24

OpenStudy (xapproachesinfinity):

\(1/3-(-7/8)=1/3+7/8\) recall \((-)(-)=+\)

OpenStudy (anonymous):

only 7/8 is negative

OpenStudy (xapproachesinfinity):

but you have minus btw the two fraction yes?

OpenStudy (anonymous):

but if it is negative, then nyour result is -13/24

OpenStudy (anonymous):

i have to show the work also, and i dont understand much about fractions in general.

OpenStudy (xapproachesinfinity):

\(\large \frac{1}{3}-\frac{-7}{8}\)? this right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

all i know so far is how to get the lcd, which is 24.

OpenStudy (xapproachesinfinity):

well think about this \(a-(-b)=?\)

OpenStudy (anonymous):

that is (1/3)--(7/8)= L.c.m =24 (8+21)/24 = 29/24

OpenStudy (anonymous):

because double minus a--b=a+b

OpenStudy (xapproachesinfinity):

from arithmetic rules \(a-(-b)=a+(-(-b))=a+b\)

OpenStudy (xapproachesinfinity):

recall that \((-)*(-)=(+)\)

OpenStudy (xapproachesinfinity):

that is also applied to rational numbers as well

OpenStudy (xapproachesinfinity):

rational numbers are basically fractions

OpenStudy (xapproachesinfinity):

So \(\large \frac{1}{3}+\frac{7}{8}=\frac{8}{24}+\frac{21}{24}=\frac{8+21}{24}\)

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

how did the 7/8 turn into 21/24?

OpenStudy (xapproachesinfinity):

well we are doing Common Denominator and that is 24 correct?

OpenStudy (anonymous):

yes.

OpenStudy (xapproachesinfinity):

we multiplied 3 by 8 to 24 and 8 by 3 to get 24 but we need to multiply top and bottom so 1/3 has to be multiplied by 8 and 7/8 must be multiplied by 3 top and bottom

OpenStudy (anonymous):

so the answer would be 29/24?

OpenStudy (xapproachesinfinity):

correct

OpenStudy (anonymous):

okay, thank you!

OpenStudy (xapproachesinfinity):

yw!